solving for x:
13.7b-6.5= -2.3b+8.3

Answer by venugopalramana(3286)

13.7b-6.5= -2.3b+8.3
13.7b+2.3b=8.3+6.5=14.8
b(13.7+2.3)=16b= 14.8
b=14.8/16 =3.7/4